已知事件 AA 和 BB 满足 P(A)=0.7P(A)=0.7 且 P(B)=0.8P(B)=0.8。证明 P(A∩B)≥0.5P(A \cap B) \ge 0.5 若 P(A′ ∪ B)=0.85P(A'\ \cup\ B)=0.85,求 P(A∣B)P(A\mid B)。说明并解释 AA 和 BB 是否为独立事件。 若 P(A∩B)=0.5P(A \cap B)=0.5,事件 A 与事件 B 之间有什么关系? 在一平面上画出 11 条互不平行的直线。若没有任意三条直线交于同一点,证明这些直线之间有 55 个交点。 若随机选择其中三个交点,求它们都在这 11 条直线中的某一条上的概率。

先不看答案,让 AI 诊断你卡在哪里?

它不会直接告诉你答案,而是问你卡在哪一步看看 AI 是怎么陪你一步步弄懂这道题的

完整解法

7((a))((i))

1. Find $P(A \cap B)$ using the addition rule.

P(A′∪B)=P(A′)+P(B)−P(A′∩B)P(A' \cup B) = P(A') + P(B) - P(A' \cap B)
0.85=(1−0.7)+0.8−P(A′∩B)0.85 = (1 - 0.7) + 0.8 - P(A' \cap B)
0.85=0.3+0.8−P(A′∩B)  ⟹  P(A′∩B)=0.250.85 = 0.3 + 0.8 - P(A' \cap B) \implies P(A' \cap B) = 0.25
Since BB consists of the mutually exclusive events (A∩B)(A \cap B) and (A′∩B)(A' \cap B):
P(B)=P(A∩B)+P(A′∩B)P(B) = P(A \cap B) + P(A' \cap B)
0.8=P(A∩B)+0.25  ⟹  P(A∩B)=0.550.8 = P(A \cap B) + 0.25 \implies P(A \cap B) = 0.55

2. Calculate the conditional probability $P(A \mid B)$.

P(A∣B)=P(A∩B)P(B)P(A \mid B) = \frac{P(A \cap B)}{P(B)}
P(A∣B)=0.550.8=1116=0.6875P(A \mid B) = \frac{0.55}{0.8} = \frac{11}{16} = 0.6875

3. Test for independence of events A and B.

For AA and BB to be independent events, P(A∩B)P(A \cap B) must equal P(A)P(B)P(A)P(B).
P(A)P(B)=0.7×0.8=0.56P(A)P(B) = 0.7 \times 0.8 = 0.56
Since P(A∩B)=0.55≠0.56P(A \cap B) = 0.55 \neq 0.56, the events AA and BB are not independent.

答案

P(A∣B)=1116P(A \mid B) = \frac{11}{16} (or 0.68750.6875); AA and BB are not independent since P(A∩B)≠P(A)P(B)P(A \cap B) \neq P(A)P(B).

7((a))((ii))

1. Calculate the probability of the union of events $A$ and $B$.

P(A∪B)=P(A)+P(B)−P(A∩B)P(A \cup B) = P(A) + P(B) - P(A \cap B)
P(A∪B)=0.7+0.8−0.5=1P(A \cup B) = 0.7 + 0.8 - 0.5 = 1

2. State the relationship between the two events based on the result.

Since P(A∪B)=1P(A \cup B) = 1, the union of events AA and BB covers the entire sample space. Therefore, AA and BB are exhaustive events.

答案

Event AA and event BB are exhaustive events.

7((b))

1. Show that the total number of intersection points is 55.

Since the 1111 lines are non-parallel and no three lines intersect at a common point, each pair of lines intersects at exactly one unique point.
Total number of intersection points=11C2\text{Total number of intersection points} = {}^{11}C_2
=11×102=55 (shown)= \frac{11 \times 10}{2} = 55 \text{ (shown)}

2. Determine the total number of ways to choose 3 points.

Total ways to choose 3extpointsfrom55=55C3\text{Total ways to choose } 3 ext{ points from } 55 = {}^{55}C_3
=55×54×533×2×1=26235= \frac{55 \times 54 \times 53}{3 \times 2 \times 1} = 26235

3. Determine the number of favourable outcomes (choosing 3 points that all lie on one of the 11 lines).

Each of the 1111 lines intersects the other 1010 lines, so there are exactly 1010 intersection points on each line.
For a specific line, the number of ways to choose 33 points on it is:
10C3=10×9×83×2×1=120{}^{10}C_3 = \frac{10 \times 9 \times 8}{3 \times 2 \times 1} = 120
Since two distinct lines share at most 11 point, no 33 points can be collinear on more than one line simultaneously. Thus, the subsets of 33 points on each of the 1111 lines are mutually exclusive.
Total favourable outcomes=11×10C3\text{Total favourable outcomes} = 11 \times {}^{10}C_3
=11×120=1320= 11 \times 120 = 1320

4. Calculate the final probability.

P(all 3 points are on one of the 11 lines)=Favourable outcomesTotal outcomesP(\text{all } 3 \text{ points are on one of the } 11 \text{ lines}) = \frac{\text{Favourable outcomes}}{\text{Total outcomes}}
=132026235= \frac{1320}{26235}
=8159= \frac{8}{159}

答案

8159\frac{8}{159}

AI 分析,可能出错

用 AI 批改,圈出写错的那一步,并自动将你的薄弱点存入错题本,形成你的错因库。

同一知识点的其他题

  1. 1已知事件 AA 和 BB 满足 P(B)=13\mathrm{P}(B) = \frac{1}{3},P(A∣B)=25\mathrm{P}(A|B) = \frac{2}{5} 且 P(A∪B)=35\mathrm{P}(A \cup B) = \frac{3}{5}。…
  2. 2凯(Kai)和里奥(Leo)两兄弟是定期训练的双项全能(跑步和骑行)运动员。在每一天中,凯骑行的概率为 34\frac{3}{4},跑步的概率为 35\frac{3}{5},两者都进行的概率为 pp。 平均而言,里奥在一周的 7 天中有…
  3. 3一对同胞 Abbey 和 Betty 还有其他六个朋友一起去参加书展。他们八个人随机排成一队等候一位著名作家的签名。求满足以下条件的概率: Abbey 排在队列第一位且 Betty 排在队列第六位, Abbey 排在队列第一位或 Betty…
  4. 4已知某种特定疾病的感染率约为万分之一(每 10 000 人中约有 1 人感染)。针对该疾病的一项检测准确率为 99.9%,即检测结果与真实情况相符的概率为 99.9%。求在已知某人检测结果呈阳性的条件下,此人确实患有该疾病的概率。 研究人员…
  5. 5对于事件 AA and BB,已知 P(A)=0.38\mathrm{P}(A) = 0.38,P(A∣B)=0.5\mathrm{P}(A \mid B) = 0.5 且 P(A∪B)=0.52\mathrm{P}(A \cup B) = 0.52。 求 P(B)\mathrm{P}(B)…